OpenStudy (anonymous):

The line passing through the points (-1,6) and (4,x) has gradient -2. Find the value of x. Show working out please.

OpenStudy (raden):

use the formula : m = (y2-y1)/(x2-x1)

OpenStudy (raden):

(x-6)/(4-(-1)) = -2 (x-6)/(4+1) = -2 (x-6)/5 = -2 solve for x = ...